OpenStudy (precal):

Optimization problem Find the point on the graph of f(x)=square root of (-x+8) so that the point (2,0) is closest to the graph.

OpenStudy (precal):

\[f(x)=\sqrt{-x+8}\]

OpenStudy (precal):

I found the derivative of f (x) but not sure what else to do with this information

OpenStudy (precal):

\[f ' (x)=\frac{ 1 }{ 2\sqrt{-x+8} }\]

OpenStudy (precal):

@ganeshie8

OpenStudy (precal):

@SolomonZelman

OpenStudy (solomonzelman):

I am not so good at these. https://www.desmos.com/calculator/ixbpyipjp7

OpenStudy (precal):

thanks

OpenStudy (solomonzelman):

What, helped ?

OpenStudy (precal):

not really

OpenStudy (solomonzelman):

Ohh, I tried... sorry-:(

OpenStudy (precal):

no problem, I have to show work using calculus

OpenStudy (precal):

@jim_thompson5910

OpenStudy (precal):

@dumbcow

OpenStudy (precal):

sorry I gotta go pick up my tadpoles, I really appreciate any help. I will read the comments later

OpenStudy (dumbcow):

ok you want to minimize distance from (2,0) to (x,f(x)) so first use distance formula \[d = \sqrt{(x-2)^2 + (8-x)} = \sqrt{x^2 -5x+12}\] its this function d(x) that you want to take derivative of \[d'(x) = \frac{2x-5}{2 \sqrt{x^2 -5x+12}} = 0\] \[x = \frac{5}{2}\]

OpenStudy (anonymous):

might i make a small suggestion?

OpenStudy (anonymous):

it is always easier to work with the square of the distance rather than the distance

OpenStudy (precal):

@dumbcow @satellite73 how does f(x) play into this?